OpenStudy (anonymous):

Solve. 54 is 24% of what?

OpenStudy (anonymous):

54 = 24% x 54 = 0.24x x = 54/0.24

OpenStudy (anonymous):

(24)/(100)*x=54 (6)/(25)*x=54 (6x)/(25)=54 6x=1350 x=225